SerCP11 2.1.P018 A race car moves such that its position fits the relationship x - (4.0 m/s)t + (0.70 m/s3)t3 where x is measured in meters and t in seconds (a) A plot of the cars position versus time is which of the following? x (m) 180 160 140 120 100 80 60 40 20 x (m) 200 180 160 140 120 100 80 60 40 20 t (s) t (s) 1.00 2.00 3.00 4.00 5.00 6.00 1.00 2.00 3.00 4.00 5.00 6.00 x (m) 25 20 15 10 x (m) 80 70. 60 50 40 30 2 1 0 t (s) 1.00 2.00 3.00 4.00 5.00 6.00 1.00 2.00 3.00 4.00 5.00 6.00 (b) Determine the instantaneous velocity of the car at t 3.4 s, using time intervals of 0.40 s, 0.20 s, and 0.10 s. (In ord

better see the limiting process keep at least three decimal places in your answer.) At 0.40 Your response differs from the correct answer by more than 10%. Double check your calculations. m/s interval from t 3.20 s to 3.60 s.) (Use the At 0.20 m/s (Use the interval from t 3.30 s to 3.50 s.) At 0.10 m/s (Use the interval from t 3.35 s to 3.45 s.) (c) Compare the average velocity during the first 3.4 s with the results of part (b). The average velocity of Your response is within 10% of the correct value. This may be due to roundoff error, or you could have a mistake in your calculation. Carry out all intermediate results to at least four-digit accuracy to minimize roundoff error. m/s is much less than the instantaneous velocity. Help? Read
Answer #1

(a)

Car's position versus time graph,

First graph is correct.

(b)

Given that,

x = 4t + 0.70t^3

Instantaneous velocity of car at t = 3.4 s,

v = dx / dt = d(4t + 0.70t^3) / dt

v = 4 + 2.1t^2

v = 4 + 2.1*(3.4)^2

v = 28.27 m/s

(c)

displacement during first 3.4 s,

x = 4t + 0.70t^3

x = 4*3.4 + 0.70*3.4^3 = 41.112

Average velocity during first 3.4 s,

vavg = 41.112 / 3.4

vavg = 12.09 m/s